The ratio if the present ages of a and b is 7:9 nine years ago the the ratio of their ages was 2:3 find their present ages

Answer by Boreal(15235)

The present ages are in a ratio of 7x:9x
nine years ago, 7x-9:9x-9 as 2:3
7x-9/(9x-9)=2/3
cross multiply
21x-27=18x-18
3x=9
x=3
The present ages are 21 and 27 ANSWER
9 years ago, they were 12 and 18, which is 2:3.